# Learning with Local Search MCMC Layers

### v1

@@ Results.
In Fig. 2, we observe that the initialization method plays an important role, and the ground-truth-based ones greatly outperform the random one. We observe that the number of Markov iterations $K$ is an important performance factor. Interestingly, the ground-truth initialization significantly improves the learning process for small $K$.
In Table 3, we compare training methods with fixed compute time budget for the layer (perturbed solver or proposed MCMC approach), which is by far the main computational bottleneck. The models are selected using a validation set and evaluated on the test set. We observe that the proposed approach significantly outperforms the perturbation-based method [6] using $\widetilde{{\bm{y}}}$ in low time limit regimes, thus allowing for faster and more efficient training. Full experimental details and additional results on the impact of temperature are given in Section B.4.
Figure 2: Test relative cost (%). Left: varying initialization method. Center: varying number of Markov iterations $K$ with random initialization. Right: varying $K$ with ground-truth initialization.
Table 3: Best test relative cost (%) for different training methods and time limits.
@@
Proposed (${\bm{y}}^{(0)}\!\!=\!{\bm{y}}$) $10.0\pm 1.7$  $12.0\pm 2.6$  $11.8\pm 2.8$  $9.1\pm 1.7$  $8.4\pm 1.7$  $7.7\pm 1.1$
Proposed (${\bm{y}}^{(0)}\!\!=\!{\bm{y}}$ + heuristic) $\mathbf{7.8}\pm 0.8$  $\mathbf{7.2}\pm 0.6$  $\mathbf{6.3}\pm 0.7$  $\mathbf{6.2}\pm 0.8$  $\mathbf{5.9}\pm 0.7$  $5.9\pm 0.6$
6 Conclusion, limitations and future work
In this paper, we introduced a principled framework for integrating NP-hard combinatorial optimization layers into neural networks without relying on exact solvers. Our approach adapts neighborhood systems from the metaheuristics community, to design structure-aware proposal distributions for combinatorial MCMC. This leads to significant training speed-ups, which are crucial for tackling larger problem instances, especially in operations research where scaling up leads to substantial value creation by reducing marginal costs. However, our framework requires opening the solver’s blackbox, and correction ratios must be implemented in the acceptance rule of the local search to guarantee convergence to the correct stationary distribution. In future work, we plan to extend our framework to large neighborhood search algorithms, which are local search heuristics that leverage neighborhood-wise exact optimization oracles.
@@ Appendix A Experiments on empirical convergence of gradients and parameters
In this section, we evaluate the proposed approach on two discrete output spaces: sets and $\kappa$-subsets. These output spaces are for instance useful for multilabel classification. We focus on these output spaces because the exact MAP and marginal inference oracles are available, allowing us to compare our gradient estimators to exact gradients.
@@ A.1 Polytopes and corresponding oracles
$\mathbb{E}_{{\pi}_{\bm{\theta},t}}\left[Y_{i}\right]$ $=\sum_{\bm{y}\in\mathcal{Y}^{d}}\frac{\exp\left(\langle\bm{\theta},\bm{y}\rangle/t\right)}{\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{d}}\exp\left(\langle\bm{\theta},\bm{y}^{\prime}\rangle/t\right)}y_{i}$
$=\sum_{\bm{y}\in\{0,1\}^{d}}\frac{\exp\left(\sum_{j=1}^{d}\theta_{j}y_{j}/t\right)}{\sum_{\bm{y}^{\prime}\in\{0,1\}^{d}}\exp\left(\sum_{j=1}^{d}\theta_{j}y^{\prime}_{j}/t\right)}y_{i}$
@@
$=\frac{0\cdot\exp(0)+1\cdot\exp\left(\theta_{i}/t\right)}{\exp(0)+\exp\left(\theta_{i}/t\right)}$
@@ A.2 Neighborhood graphs
That is, two vertices are neighbors if their Hamming distance is at most $r$. This graph is regular, with degree $|\mathcal{N}_{\leq}^{r}(\bm{y})|=\sum_{i=1}^{r}\binom{d}{i}$. This graph is naturally connected, as any binary vector $\bm{y}^{\prime}$ can be reached from any other binary vector $\bm{y}$ in $||\bm{y}^{\prime}-\bm{y}||_{1}$ moves, by flipping each bit where $y_{i}^{\prime}\neq y_{i}$, iteratively. Indeed, this trajectory consists in moves between vertices with Hamming distance equal to $1$, and are therefore along edges of the neighborhood graph, regardless of the value of $r$.
@@
These graphs, on the contrary, are not always connected: indeed, if $r$ is even, they contain two connected components (binary vectors with an even sum, and binary vectors with an odd sum). We only use such graphs when experimenting with neighborhood mixtures (see Algorithm 2), by aggregating them into a connected graph.
@@
That is, two vertices are neighbors if one can be reached from the other by performing $s$ “swaps”, each swap corresponding to flipping a $1$ to a 00 and vice-versa. This ensures that the resulting vector is still in $\mathcal{Y}_{\kappa}^{d}$. All $s$ swaps must be performed on distinct components. The resulting graph is known as the Generalized Johnson Graph $J(d,\kappa,\kappa-s)$, or Uniform Subset Graph [12]. It is a regular graph, with degree $|\mathcal{N}^{s}(\bm{y})|=\binom{\kappa}{s}\binom{d-\kappa}{s}$. It is proved to be connected in Jones [24], except if $d=2\kappa$ and $s=\kappa$ (in this case, it consists in $\frac{1}{2}\binom{d}{\kappa}$ disjoint edges).
When $s=1$, the neighborhood graph is the Johnson Graph $J(d,\kappa)$, which coincides with the graph associated to the polytope $\operatorname{\mathsf{conv}}(\mathcal{Y}_{\kappa}^{d})=\Delta_{d,\kappa}$ [43].
@@ A.3 Convergence to exact gradients
where $\bm{y}^{(k)}$ is the $k$-th iterate of Algorithm 1 with maximization direction $\bm{\theta}$, final temperature $t$, and $K_{0}$ is a number of burn-in (or warm-up) iterations. The obtained estimator is compared to the exact expectation by performing marginal inference as described in Section A.1 (with a closed-form formula in the case of $\mathcal{Y}^{d}$, and by dynamic programming in the case of $\mathcal{Y}_{\kappa}^{d}$).
@@
$\frac{1}{M}\sum_{i=1}^{M}||\mathbb{E}_{{\pi}_{\bm{\Theta}_{i}},t}\left[Y\right]-\tilde{\mathbb{E}}(\bm{\Theta}_{i},T)||_{2}^{2},$
@@
First, we evaluate the impact of $K_{0}$, the number of burn-in iterations. We use a truncated geometric cooling schedule $t_{k}=\max(\gamma^{k}\cdot t_{0},\,t)$ with $\gamma=0.995$. The initial temperature $t_{0}$ is set to $1/(\gamma^{K_{0}})$, so that $\forall k\geq K_{0}+1,\,t_{k}=t=1$. The results are gathered in Fig. 3.
@@
We then evaluate the impact of the final temperature $t$ on the difficulty of the estimation task (different temperatures lead to different target expectations). The results for the small scale are gathered in Fig. 4, and the results for the large scale are gathered in Fig. 5.
@@
Finally, we evaluate the impact of the number of parallel Markov chains $C$ on the estimation. The results are gathered in Fig. 6.
@@
(a) SA, hypercube
(b) SA, top-$\kappa$ polytope
@@ A.4 Convergence to exact parameters
In this section, we conduct experiments in the unsupervised setting described in Section 4.4. As a reminder, the empirical $L_{N}$ and population $L_{{\bm{\theta}}_{0}}$ Fenchel-Young losses are given by:
@@
[…]
Figure 7: Comparison of stochastic gradient trajectories for a SA / M-H oracle on $\mathcal{Y}^{2}$ and unbiased stochastic gradients obtained via Monte Carlo sampling. Increasing the number of Markov chain iterations yields smoother trajectories, similar to the effect of using more Monte Carlo samples in the case of perturbation-based methods [6].
@@
We proceed by first randomly generating true parameters $\bm{\Theta}_{0}\in\mathbb{R}^{M\times d}$, with $M$ being a number of problem instances we average on (in order to reduce noise in our observations), by sampling $\bm{\Theta}_{i,j}\sim\mathcal{N}(0,1)$ independently. The goal is to learn each parameter vector $(\bm{\Theta}_{0})_{i}\in\mathbb{R}^{d},i\in[M]$, as $M$ independent problems. The model is randomly initialized at $\hat{\bm{\Theta}}_{0}$, and updated with Adam [25] to minimize the loss. In order to better separate noise due to the optimization process and noise due to the sampling process, we use the population loss $L_{(\bm{\Theta}_{0})_{i}}$ for general experiments, and use the empirical loss $L_{N}$ only when focusing on the impact of the dataset size $N$. In this case, we create a dataset $Y\in\mathbb{R}^{M\times N\times d}$, with $N$ being the number of samples, by sampling independently $Y_{i,j}\sim{\pi}_{(\bm{\Theta}_{0})_{i}},\;\forall i\in[M],\,\forall j\in[N]$.
@@
    $N$, the number of samples in the dataset.
@@
In both cases, the best loss value that can be reached is positive but cannot be computed: it corresponds to the constants $C_{1}$ and $C_{2}$ in Eq. 8 and Eq. 9. Thus, we also provide "stretched" figures, where we plot the loss minus the best loss found during the optimization process.
@@
$\frac{1}{M}\sum_{i=1}^{M}||(\bm{\Theta}_{0})_{i}-(\hat{\bm{\Theta}}_{n})_{i}||_{2}^{2}.$
As the top-$\kappa$ polytope is of dimension $d-1$, the model is only specified up to vectors orthogonal to the direction of the smallest affine subspace it spans. Thus, in this case, we measure instead:
$\frac{1}{M}\sum_{i=1}^{M}||\operatorname*{\mathsf{P}^{\perp}}_{D}\left((\bm{\Theta}_{0})_{i}\right)-\operatorname*{\mathsf{P}^{\perp}}_{D}\left((\hat{\bm{\Theta}}_{n})_{i}\right)||_{2}^{2},$
@@
First, we evaluate the impact of $K$, the number of inner iterations, i.e., the length of each Markov chain. The results are gathered in Fig. 8.
@@
We now evaluate the impact of the number of Markov chains $C$ run in parallel to perform each gradient estimation on the learning process. The results are gathered in Fig. 9.
@@
Then, we evaluate the impact of the method to initialize each Markov chain used for gradient estimation. The persistent method consists in setting $\bm{y}^{(n+1,0)}=\bm{y}^{(n,K)}$, the data-based method consists in setting $\bm{y}^{(n+1,0)}={\bm{y}}_{i}$ with $i\sim\mathcal{U}([N])$, and the random method consists in setting $\bm{y}^{(n+1,0)}\sim\mathcal{U}(\mathcal{Y})$ (see Section B.2 and Table 4 for a detailed explanation). The results are gathered in Fig. 10.
@@
We now evaluate the impact of the number of samples $N$ from $\pi_{\bm{\theta}_{0}}$ (i.e., the size of the dataset $\left({\bm{y}}_{i}\right)_{i=1}^{N}$) on the estimation of the true parameter $\bm{\theta}_{0}$. The results are gathered in Fig. 11.
@@
Finally, we evaluate the impact of the use of neighborhood mixtures. To do so, we use mixtures $\{\mathcal{N}_{=}^{r_{s}}\}_{s=1}^{S}$, once with $\{r_{s}\}_{s=1}^{S}=\{5\}$ opposed to $\{r_{s}\}_{s=1}^{S}=\{1,5\}$, and once with $\{r_{s}\}_{s=1}^{S}=\{6\}$ (which gives a reducible Markov chain as $6$ is even, so that the individual neighborhood graph $\mathcal{N}_{=}^{6}$ is not connected, and has to connected components) opposed to $\{r_{s}\}_{s=1}^{S}=\{1,2,3,6\}$. The results are gathered in Fig. 12.

### v2

@@ Results.
We use the competition’s metric: the routing cost over full instances with multiple dispatching waves, relative to the anticipative oracle $f^{\star}\!$. In Fig.˜2, initializing the Markov chain with the ground-truth solution clearly outperforms a random start (even more so when refined by the fast initialization heuristic used by $\tilde{{\bm{y}}}$), and performance increases with the MCMC iteration number $K$.
In Table˜3, we compare training methods under a fixed time budget for the layer’s forward pass (the main computational bottleneck). We observe that our approach significantly outperforms the perturbation-based method in low time-limit regimes (11-$100$ ms), thus enabling faster and more efficient training. Full experimental details and additional results are in Section˜C.7.
Figure 2: Test relative cost (%) w.r.t. $f^{\star}$. Left: varying initialization method. Center: varying number of Markov iterations $K$, random initialization. Right: varying $K$, ground-truth initialization.
Table 3: Best test relative cost (%) w.r.t. $f^{\star}$ for different training methods and time limits.
@@
Proposed (${\bm{y}}^{(0)}\!\!=\!{\bm{y}}$) $10.0\pm 1.7$  $2.0\pm 2.6  11$$1.8\pm 2.8$  $9.1\pm 1.7$  $8.4\pm 1.7$  $7.7\pm 1.1$
Proposed (${\bm{y}}^{(0)}\!\!=\!\text{heuristic}({\bm{y}})$) $\mathbf{7.8}\pm 0.8$  $\mathbf{7.2}\pm 0.6$  $\mathbf{6.3}\pm 0.7$  $\mathbf{6.2}\pm 0.8$  $\mathbf{5.9}\pm 0.7$  $5.9\pm 0.6$
5.2 Learning to predict binary vectors
Setup.
To further validate the proposed gradient estimators, we use a synthetic unconditional learning task with hypercube output space, ${\mathcal{Y}}=\{0,1\}^{d}$. This setting is ideal for controlled experiments because the Gibbs distribution $\pi_{{\bm{\theta}},t}$ is fully factorized, leading to trivial sampling and a tractable closed-form expectation ${\mathbb{E}}_{\pi_{{\bm{\theta}},t}}[Y]=\sigma({\bm{\theta}}/t)$, where $\sigma$ is the logistic sigmoid function. This allows us to both faithfully generate datasets from a known distribution $\pi_{{\bm{\theta}},t}$ , and to minimize the population Fenchel-Young loss $L_{{\bm{\theta}}}$ directly (see Section˜4.4 for its definition). The latter lets us decouple the noise from our MCMC estimator from the statistical noise inherent in finite datasets.
In all experiments, the goal is to recover a known “true” parameter vector ${\bm{\theta}}_{0}$ from independent samples $({\bm{y}}_{i})_{i=1}^{N}\sim(\pi_{{\bm{\theta}}_{0},t})^{\otimes N}$. We summarize our key findings in Fig.˜3, which shows the distance to ${\bm{\theta}}_{0}$ along a stochastic gradient trajectory, either minimizing $L_{N}$ (left) or $L_{{\bm{\theta}}_{0}}$ (center, right). Full experimental and theoretical details are available in Appendix˜A, together with additional results on both the hypercube and the top-$\kappa$ polytope.
Results.
The results highlight three important aspects for effective learning. First, persistent and data-based initializations for the MCMC chains are critical (see Section˜B.3 for a detailed discussion), vastly outperforming random restarts, which introduce systematic bias in the gradient estimation (Fig.˜3, center). Second, a larger dataset size $N$ provides a better approximation of the population loss, leading to a more accurate parameter recovery (Fig.˜3, left), in line with Section˜4.4. Finally (defining Hamming distance-based neighborhood systems $(\mathcal{N}_{r_{s}})_{s=1}^{S}$ by ${\bm{y}}^{\prime}\in\mathcal{N}_{r_{s}}(\bm{y})\Leftrightarrow d_{H}\left(\bm{y},\,\bm{y}^{\prime}\right)=r_{s}$), using a mixture of proposals with Algorithm˜2 (e.g., with $r_{s}\in\{1,2,3,6\}$) enables more effective exploration, improving convergence compared to a single proposal type (Fig.˜3, right).
(a) Varying initialization
(b) Varying dataset size NN
(c) Varying neighborhoods
Figure 3: Squared distance $\hat{{\bm{\theta}}}_{n}-{\bm{\theta}}_{0}||_{2}^{2}$ to the true parameter over optimization steps.
6 Conclusion
In this paper, we introduced a principled framework for integrating NP-hard combinatorial optimization layers into neural networks without relying on exact solvers. Our approach adapts neighborhood systems from the metaheuristics community, to design structure-aware proposal distributions for combinatorial MCMC. This leads to significant training speed-ups, enabling to tackle larger problem instances, which is crucial in operations research, where scaling up leads to substantial value creation by reducing marginal costs. In future work, we plan to extend our framework to large neighborhood search algorithms, which are heuristics that leverage neighborhood-wise exact optimization oracles.
@@ Appendix A Experiments on empirical convergence of gradients and parameters
In this section, we evaluate the proposed approach on two discrete output spaces: sets and $\kappa$-subsets. These output spaces are for instance useful for multilabel classification. We focus on these output spaces because the exact MAP and marginal inference oracles are available, allowing us to compare our gradient estimators to exact gradients. We set $\varphi\equiv 0$ in these experiments.
@@ A.1 Polytopes and corresponding oracles
$\mathbb{E}_{{\pi}_{\bm{\theta},t}}\left[Y_{i}\right]$ $=\sum_{\bm{y}\in\mathcal{Y}^{d}}\frac{\exp\left(\langle\bm{\theta},\bm{y}\rangle/t\right)}{\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{d}}\exp\left(\langle\bm{\theta},\bm{y}^{\prime}\rangle/t\right)}y_{i}=\sum_{\bm{y}\in\{0,1\}^{d}}\frac{\exp\left(\sum_{j=1}^{d}\theta_{j}y_{j}/t\right)}{\sum_{\bm{y}^{\prime}\in\{0,1\}^{d}}\exp\left(\sum_{j=1}^{d}\theta_{j}y^{\prime}_{j}/t\right)}y_{i}$
@@ A.2 Neighborhood graphs
That is, two vertices are neighbors if their Hamming distance is at most $r$. This graph is regular, with degree $|\mathcal{N}_{\leq}^{r}(\bm{y})|=\sum_{i=1}^{r}\binom{d}{i}$. This graph is naturally connected, as any binary vector $\bm{y}^{\prime}$ can be reached from any other binary vector $\bm{y}$ in $\bm{y}^{\prime}-\bm{y}||_{1}$ moves, by flipping each bit where $y_{i}^{\prime}\neq y_{i}$, iteratively. Indeed, this trajectory consists in moves between vertices with Hamming distance equal to $1$, and are therefore along edges of the neighborhood graph, regardless of the value of $r$.
@@
These graphs, on the contrary, are not always connected: indeed, if $r$ is even, they contain two connected components (binary vectors with an even sum, and binary vectors with an odd sum). We only use such graphs when experimenting with neighborhood mixtures (see Algorithm˜2), by aggregating them into a connected graph.
@@
That is, two vertices are neighbors if one can be reached from the other by performing $s$ “swaps”, each swap corresponding to flipping a $1$ to a 0 and vice-versa. This ensures that the resulting vector is still in $\mathcal{Y}_{\kappa}^{d}$. All $s$ swaps must be performed on distinct components. The resulting graph is known as the Generalized Johnson Graph $J(d,\kappa,\kappa-s)$, or Uniform Subset Graph [Chen and Lih, 1987]. It is a regular graph, with degree $|\mathcal{N}^{s}(\bm{y})|=\binom{\kappa}{s}\binom{d-\kappa}{s}$. It is proved to be connected in Jones [2005], except if $d=2\kappa$ and $s=\kappa$ (in this case, it consists in $\frac{1}{2}\binom{d}{\kappa}$ disjoint edges).
When $s=1$, the neighborhood graph is the Johnson Graph $J(d,\kappa)$, which coincides with the graph associated to the polytope $\operatorname{\mathsf{conv}}(\mathcal{Y}_{\kappa}^{d})=\Delta_{d,\kappa}$ [Rispoli, 2008].
@@ A.3 Convergence to exact gradients
where $\bm{y}^{(k)}$ is the $k$-th iterate of Algorithm˜1 with maximization direction $\bm{\theta}$, final temperature $t$, and $K_{0}$ is a number of burn-in (or warm-up) iterations. The obtained estimator is compared to the exact expectation by performing marginal inference as described in Section˜A.1 (with a closed-form formula in the case of $\mathcal{Y}^{d}$, and by dynamic programming in the case of $\mathcal{Y}_{\kappa}^{d}$).
@@
[…]
@@
First, we evaluate the impact of $K_{0}$, the number of burn-in iterations. We use a truncated geometric cooling schedule $t_{k}=\max(\gamma^{k}\cdot t_{0},\,t)$ with $\gamma=0.995$. The initial temperature $t_{0}$ is set to $1/(\gamma^{K_{0}})$, so that $\forall k\geq K_{0}+1,\,t_{k}=t=1$. The results are gathered in Fig.˜4.
@@
We then evaluate the impact of the final temperature $t$ on the difficulty of the estimation task (different temperatures lead to different target expectations). The results for the small scale are gathered in Fig.˜5, and the results for the large scale are gathered in Fig.˜6.
@@
Finally, we evaluate the impact of the number of parallel Markov chains $C$ on the estimation. The results are gathered in Fig.˜7.
@@
(a) Hypercube
(b) Top-$\kappa$ polytope
@@ A.4 Convergence to exact parameters
In this section, we conduct experiments in the unconditional setting described in Section˜4.4. As a reminder, the empirical $L_{N}$ and population $L_{{\bm{\theta}}_{0}}$ Fenchel-Young losses are given by:
@@
As an introductory example, we display stochastic gradient trajectories in Fig.˜8. The parameter $\bm{\theta}\in\mathbb{R}^{d}$ is updated following Eq.˜7 to minimize the population loss $L_{\bm{\theta}_{0}}$ defined in Eq.˜10, with $\bm{\theta}_{0}=(1/2,\,1/2)$. The polytope used is the $2$-dimensional hypercube $\mathcal{Y}^{2}$, with neighborhood graph $\mathcal{N}_{1}$ (neighbors are adjacent vertices of the square). We present trajectories obtained using MCMC-sampled gradients, comparing results from both $1$ and $100$ Markov chain iterations with Algorithm˜1. For comparison, we include trajectories obtained using Monte Carlo-sampled (i.e., unbiased) gradients, using 1 and 100 samples.
Figure 8: Comparison of stochastic gradient trajectories for a SA / M-H oracle on $\mathcal{Y}^{2}$ and unbiased stochastic gradients obtained via Monte Carlo sampling. Increasing the number of Markov chain iterations yields smoother trajectories, similar to the effect of using more Monte Carlo samples in the case of perturbation-based methods [Berthet et al., 2020].
@@
We proceed by first randomly generating true parameters $\bm{\Theta}_{0}\in\mathbb{R}^{M\times d}$, with $M$ being a number of problem instances we average on (in order to reduce noise in our observations), by sampling $\bm{\Theta}_{i,j}\sim\mathcal{N}(0,1)$ independently. The goal is to learn each parameter vector $(\bm{\Theta}_{0})_{i}\in\mathbb{R}^{d},i\in[M]$, as $M$ independent problems. The model is randomly initialized at $\hat{\bm{\Theta}}_{0}$, and updated with Adam [Kingma and Ba, 2017] to minimize the loss. In order to better separate noise due to the optimization process and noise due to the sampling process, we use the population loss $L_{(\bm{\Theta}_{0})_{i}}$ for general experiments, and use the empirical loss $L_{N}$ only when focusing on the impact of the dataset size $N$. In this case, we create a dataset $Y\in\mathbb{R}^{M\times N\times d}$, with $N$ being the number of samples, by sampling independently $Y_{i,j}\sim{\pi}_{(\bm{\Theta}_{0})_{i}},\;\forall i\in[M],\,\forall j\in[N]$.
@@
    NN, the number of samples in the dataset.
@@
In both cases, the best loss value that can be reached is positive but cannot be computed: it corresponds to the constants $C_{1}$ and $C_{2}$ in Eq.˜9 and Eq.˜10. Thus, we also provide "stretched" figures, where we plot the loss minus the best loss found during the optimization process.
@@
[…]
As the top-$\kappa$ polytope is of dimension d1d-1, the model is only specified up to vectors orthogonal to the direction of the smallest affine subspace it spans. Thus, in this case, we measure instead:
[…]
@@
First, we evaluate the impact of $K$, the number of inner iterations, i.e., the length of each Markov chain. The results are gathered in Fig.˜9.
@@
We now evaluate the impact of the number of Markov chains $C$ run in parallel to perform each gradient estimation on the learning process. The results are gathered in Fig.˜10.
@@
Then, we evaluate the impact of the method to initialize each Markov chain used for gradient estimation. The persistent method consists in setting $\bm{y}^{(n+1,0)}=\bm{y}^{(n,K)}$, the data-based method consists in setting $\bm{y}^{(n+1,0)}={\bm{y}}_{i}$ with $i\sim\mathcal{U}([N])$, and the random method consists in setting $\bm{y}^{(n+1,0)}\sim\mathcal{U}(\mathcal{Y})$ (see Section˜B.3 and Table˜4 for a detailed explanation). The results are gathered in Fig.˜11.
@@
We now evaluate the impact of the number of samples $N$ from $\pi_{\bm{\theta}_{0}}$ (i.e., the size of the dataset $\left({\bm{y}}_{i}\right)_{i=1}^{N}$) on the estimation of the true parameter $\bm{\theta}_{0}$. The results are gathered in Fig.˜12.
@@
Finally, we evaluate the impact of the use of neighborhood mixtures. To do so, we use mixtures $\{\mathcal{N}_{=}^{r_{s}}\}_{s=1}^{S}$, once with $\{r_{s}\}_{s=1}^{S}=\{5\}$ opposed to $\{r_{s}\}_{s=1}^{S}=\{1,5\}$, and once with $\{r_{s}\}_{s=1}^{S}=\{6\}$ (which gives a reducible Markov chain as $6$ is even, so that the individual neighborhood graph $\mathcal{N}_{=}^{6}$ is not connected, and has to connected components) opposed to $\{r_{s}\}_{s=1}^{S}=\{1,2,3,6\}$. The results are gathered in Fig.˜13.